# Romance, Relief, and Regret: Teen Narratives of Chatbot Overreliance

### v2

@@ Discussion
Applying this to GenAI means shifting from censorship-focused models to ones that promote reflection, co-regulation, and guided disengagement. Platforms could offer usage tracking, emotional check-ins, or personalized limits—like impulse-control tools from online safety research. Participatory design features, like customizable moderation settings or teen advisory councils (Yoon, Zhang, and Seering 2025), would give teens more control over how boundaries are set. Rather than sudden content removal, moderation systems could use transitional messages, explanations, or alternatives to help teens process changes. These strategies treat moderation not as a constraint, but as a developmental tool that fosters emotional growth, builds digital resilience, and supports teens’ existing strategies for navigating risk.
Limitations and Future Work
@@
While our study provided insights into overreliance characteristics on Character.AI, it centered on a single platform, which limits the generalizability of our findings to other AI companions with different designs, policies, or user communities. Future research should compare multiple platforms to examine how various features, moderation practices, and interface designs contribute to patterns of overuse. Additionally, our dataset consisted of voluntarily posted Reddit content, which may overrepresent individuals who realized their overreliance and tend to share their experiences online. Those who are less reflective or not active online may not post publicly, resulting in potential self-selection bias. Future researchers should consider mixed-method approaches such as surveys or structured interviews that capture a wider range of experiences, including those who use chatbots casually or positively. Also, to answer RQ2, we analyzed our data by mapping it to a behavioral addiction model, which may have oversimplified the complex emotional dynamics at play. Future research should consider alternative frameworks such as attachment (Xie, Pentina, and Hancock 2023) or parasocial theory (Horton and Richard Wohl 1956) for more insights. Also, the lack of verified demographic data may have limited our ability to examine how factors like age, gender, or background shape chatbot overreliance, suggesting a need for future studies with direct participant engagement. Lastly, the posts analyzed were in English, potentially excluding cultural differences in how teens understand or engage with AI companions, which suggests the importance of future cross-cultural research to understand global teen experiences with chatbots.
Conclusion
We found a growing tendency among adolescents to seek emotional support and entertainment through AI-driven chatbots while getting overly attached to them. The design of supportive AI systems should be informed by developmental psychology and human-centered design principles that serves developmental needs of adolescent and safeguard their well-being in such critical stage. We advocate for a proactive approach in which AI chatbot interfaces are designed not only to respond empathetically but also to promote healthy coping strategies, encourage critical thinking, and facilitate pathways to human support when appropriate. By doing so, these technologies can serve as meaningful, developmentally appropriate tools that complement—rather than replace—existing support systems in adolescents’ lives.
Ethical Considerations
This study analyzed publicly available Reddit posts self-disclosed as 13-17 years olds discussing their use of Character.AI. Given the public nature of the data and the absence of interaction with human subjects, this study was deemed as not a human subject research by our Institutional Review Board (IRB). We followed standard practices for ethical data use by anonymizing and paraphrasing excerpts to reduce potential harm or re-identification. We approached these posts with sensitivity, especially given their deeply personal and emotional nature. While the posts were historical and we did not encounter content that required immediate intervention (e.g., disclosures of self-harm or violence), we recognize that many posts reflected psychological vulnerability. We aim to ensure our findings support teen well-being rather than stigmatize their behavior or overuse.
Researchers Positionality
All authors come from a technical background in computer science and human-computer interaction (HCI), with expertise in AI and NLP. Our training does not include formal education in psychology, which influenced our choice to adopt and apply an established psychological model of behavioral addiction without adaptation. This decision reflects our respect for existing disciplinary frameworks and our intent not to reinterpret clinical or behavioral concepts beyond our expertise. We approach this work from a technical and design perspective, aiming to better understand how AI systems affect real-world behaviors—especially among young users. Our position is that responsible HCI research must critically engage with the implications of AI technologies, even when coming from a design or systems lens. Our interest is not in pathologizing teen behavior, but in informing safer, accountable design.
Impact Statement
Our intension for this study is to contribute to a more grounded understanding of how teens use emotionally responsive chatbots and the risks of overreliance. We believe our work can help shift attention toward designing AI companions that support youth development rather than inadvertently causing distress. Our findings offer a foundation for further research—especially studies from psychology, education, and public health—that could develop interventions or tools to support teens in regulating their chatbot use.
This work may also inform designers and developers of AI systems, highlighting how platform affordances shape user dependence and emotional attachment. While we mention Character.AI due to the scope of our data, our critique is not directed at any specific company. The patterns we identify are likely to emerge across many platforms offering emotionally intelligent AI. We caution against using our results to single out one provider or product and instead advocate for a broader discourse on platform responsibility, teen vulnerability, and inclusive safety design. We also recognize a potential risk: that the study might be misused to further sensationalize or moralize teen chatbot use. Our intent is not to stigmatize teens or their emotional needs, but to create space for thoughtful design and policy responses to a growing and under-examined phenomenon.

### v3

@@ 5.4. Ease of Exit: Design for Disengagement
Beyond conversational prompts, platforms could integrate usage tracking, emotional check-ins, or personalized limits such as impulse-control tools (Wisniewski et al., 2017). Participatory design features like customizable moderation settings or teen advisory councils (Yoon et al., 2025) would further give teens a voice in shaping their digital environment. Instead of sudden content removal, moderation systems could use transitional messages, explanations, or alternatives to help teens process changes. These strategies frame moderation not as a constraint, but as a developmental tool that fosters emotional growth, builds digital resilience, and supports teens’ own approaches to navigating risk.
Unlike gaming or social media, where overuse is often linked to competition, reward cycles, or social validation (Wang and Wang, 2025; Kaya et al., 2024), conversational AI companions provide personalized, reciprocal, and emotionally immersive experiences. Teens in our study did not simply consume or compete; they engaged in ongoing, two-way dialogues with AI characters that could recall past conversations, adapt to their emotional states, and simulate intimacy. These qualities, such as personalization, multimodality, and memory, set AI companions apart from earlier technologies and make overreliance harder to disentangle from authentic-feeling relationships. This underscores the need for further research on the unique characteristics of these relationships and how challenges specific to companion chatbots should be addressed.
5.5. Limitations and Future Work
@@
While our study provided insights into overreliance characteristics on Character.AI, it centered on a single platform, which limits the generalizability of our findings to other AI companions with different designs, policies, or user communities. Future research should compare multiple platforms to examine how various features, moderation practices, and interface designs contribute to patterns of overuse. Additionally, our dataset consisted of voluntarily posted Reddit content, which may overrepresent individuals who realized their overreliance and tend to share their experiences online. Those who are less reflective or not active online may not post publicly, resulting in potential self-selection bias. Future researchers should consider mixed-method approaches such as surveys or structured interviews that capture a wider range of experiences, including those who use chatbots casually or positively. Also, to answer RQ2, we analyzed our data by mapping it to a behavioral addiction model, which may have oversimplified the complex emotional dynamics at play. We note that future work should compare multiple frameworks, including alternative perspectives such as attachment (Xie et al., 2023) or parasocial theory (Horton and Richard Wohl, 1956b), to test the robustness of different interpretations. Also, the lack of verified demographic data may have limited our ability to examine how factors like age, gender, or background shape chatbot overreliance, suggesting a need for future studies with direct participant engagement. Lastly, since all authors were fluent in English, the posts analyzed were in this language, potentially excluding cultural differences in how teens understand or engage with AI companions, which suggests the importance of future cross-cultural research to understand global teen experiences with chatbots. While our dataset does not represent teens globally, it reflects a segment of English-speaking adolescents who are active online and engaging with AI companions. This linguistic boundary is common in qualitative internet research and does not invalidate the themes we observed.
6. Conclusion
This study shows that teens often begin using AI chatbots for comfort or entertainment, but many gradually become more attached and reliant. Their experiences reflect features of behavioral addiction, with consequences that include disrupted sleep, academic struggles, and strained relationships. At the same time, some teens disengage after recognizing harm, reconnecting offline, or encountering platform restrictions. By mapping these patterns to an online behavioral addiction framework, the study provides one of the first teen-centered accounts of AI companion overreliance. It highlights the unique role of emotional attachment and simulated intimacy in shaping teen experiences and introduces the CARE design framework as a practical guide for healthier interactions. Looking ahead, future research should compare platforms, use broader methods to capture diverse user experiences, and test design strategies that support constructive disengagement. Recent advances in AI create risks we have never faced before, yet they also open new opportunities. Designers now carry the responsibility to build systems with empathy, nuance, and attention to detail to not only protect teens from harm but also help them cultivate resilience, growth, and greater fulfillment in their lives.